how do you solve this...
4x + 3(2x + 3)=
can you please show me the working out? thanks Lauren.

Answer by stanbon(75887)

4x + 3(2x + 3)=
-----
4x + 3*2x + 3*3 =
-----
4x + 6x + 9 =
------
10x + 9
================
Cheers,
Stan H.